### 11 Jan 2024 v1

# In situ coherent X-ray scattering reveals polycrystalline structure and discrete annealing events in strongly-coupled nanocrystal superlattices

Abstract
Solution-phase bottom up self-assembly of nanocrystals into superstructures such as ordered superlattices is an attractive strategy to generate functional materials of increasing complexity, including very recent advances that incorporate strong interparticle electronic coupling. While the self-assembly kinetics in these systems have been elucidated and related to the product characteristics, the weak interparticle bonding interactions suggest the superstructures formed could continue to order within the solution long after the primary nucleation and growth have occurred, even though the mechanism of annealing remains to be elucidated. Here, we use a combination of Bragg coherent diffractive imaging and X-ray photon correlation spectroscopy to create real-space maps of supercrystalline order along with a real-time view of the strain fluctuations in aging strongly coupled nanocrystal superlattices while they remain suspended and immobilized in solution. By combining the results, we deduce that the self-assembled superstructures are polycrystalline, initially comprising multiple nucleation sites, and that shear avalanches at grain boundaries continue to increase crystallinity long after growth has substantially slowed. This multimodal approach should be generalizable to characterize a breadth of materials in situ in their native chemical environments, thus extending the reach of high-resolution coherent X-ray characterization to the benefit of a much wider range of physical systems.
I Introduction
Recently, a new class of bottom-up self-assembled nanomaterials has been developed that achieves a combination of long-range order and electrical conductivity by leveraging electrostatic, rather than steric, colloidal stabilization of individual nanocrystal (NC) building blocks[1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. Electrostatic colloidal stabilization is obtained by binding compact, inorganic multivalent metallochalcogen ligands to the NC surface[1]. We previously studied the solution-phase assembly kinetics of such NCs in situ with small-angle X-ray scattering (SAXS) and found fast (sub-ms) condensation of NCs into a superlattice (SL), a crystal of NCs, upon quenching via increasing the solution ionic strength[13]. In situ SAXS also shows that SL annealing progresses in solution at a small power law rate, but the nature of the coarsening events remains to be determined. Furthermore, the size and order of the resulting SL structures are seen to depend on the solution ionic strength, or quench depth, but the shape of the grains and nature of the observed heterogeneities during the annealing process are unknown, especially given the unusual interparticle interactions at play in this system[1].
To assess the structure and heterogeneity of SLs during annealing and the nature of the annealing process itself requires direct space visualization of a three-dimensional (3D) object in solution at the nanoscale and characterization of similarly small fluctuations in the structure as a function of time. Coherent hard X-ray scattering[14] techniques such as coherent diffractive imaging (CDI)[15, 16, 17, 18, 19, 20, 21, 22, 23], X-ray photon correlation spectroscopy (XPCS)[24, 25, 26, 27, 28, 29, 30, 31, 32], or a combination of the two[33] in principle allow such characterizations. In this particular case, however, they also require a high degree of translational and rotational immobilization of individual condensed structures in a focused X-ray beam in solution for observation over at least tens of minutes.
Here, we capture a snapshot of an evolving SL formed from electrostatically stabilized NCs and also observe its annealing events to elucidate their nature. To achieve stability in solution for Bragg diffraction and CDI reconstruction, we developed a strategy to immobilize SLs in capillaries filled with quartz wool. Performing Bragg CDI[34, 35, 36, 37] on the fcc (111) SL peak, we resolve a 600 nm core of apparent high electron density, surrounded by a few-micron apparently lower electron density region. We attribute the lower densities to crystallites that are not well-aligned with the measured (111) peak. Moreover, in using XPCS on the same SL, we are able to both validate the sufficient degree of immobilization in solution and elucidate the nature of SL annealing. Taken together, we conclude that SLs in solution form via a fast nucleation and growth step resulting in structures with a high density of grain boundaries. Subsequent annealing of these grain boundaries via discrete shearing events enables further ordering of the SLs.
II Results
To perform the experiments, SLs of 4.6-nm diameter Au NCs capped with thiostannate () ligands are prepared by adding 75 mM salt to an electrostatically stabilized colloidal solution in hydrazine (), as described previously[1, 13] (Fig. 1(a)). A typical small-angle X-ray scattering (SAXS) intensity pattern, , as a function of momentum transfer, , of these samples measured previously is shown in Fig. 1(b). Approximately 24 hours after the SLs are formed, they are injected into a thin-walled borosilicate capillary into which quartz wool has been inserted to immobilize them while maintaining their solvent environment to enable further evolution (Fig. 1(c), S1). The capillaries are securely sealed while still in an inert environment, and are mounted and moved into the center of rotation at the X-ray correlation spectroscopy (XCS) endstation of the Linac Coherent Light Source (LCLS) [38] (Fig. 1(d)). The experiment was carried out at 11.8 keV, just below the Au absorption edge, using the Si(111) monochromator in a SAXS geometry with a sample-to-detector distance of 7.7 m, measuring to 0.15 in order to capture the (111) SL peak as shown in Fig. 1(e). The X-ray beam, consisting of 50 fs pulses at 120 Hz, was focused with low divergence to m and attenuated to 5 photon/s to preserve the integrity of the sample and its solution environment. To characterize aspects of the data we commonly refer below to the location in momentum space on the detector relative to the direct beam () and also using a reduced scattering vector , defined as , where is the (111) reciprocal lattice vector.
Coherent diffractive imaging. To elucidate the potentially heterogeneous structure and morphology of NC SLs, a full reciprocal space sampling of the SL (111) peak was collected. The sample was rotated over a 2 range in a total of 128 steps once a SL peak was identified on the detector. After suitable pre-processing, we then performed an error reduction/hybrid input-output phase retrieval algorithm [39, 34] on the three-dimensional SL (111) peak (see Methods), with a representative detector image at a single angular step shown in 2(a). The initial object support (the region outside of which all density values are set to zero) was taken as a sphere of 1.5 m radius and enforced to be roughly spherical throughout the reconstruction process.
2(b) presents the apparent three-dimensional real space electron density, replotted in an orthonormal coordinate basis (see supplement, Fig. S2). Each (cubic) voxel has a side length of 0.071 m. The real space resolution is given by divided by the dimensions of the sampled reciprocal space region and is then nearest-neighbor interpolated to obtain isotropic sampling. The corresponding reconstructed diffraction pattern, obtained via Fourier transform of the electron density, is shown in 2(c). Comparison between 2(a) and 2(c) shows good agreement between the measured and reconstructed diffraction patterns, especially in brighter regions where the signal-to-noise ratio in the data is higher. In particular, the reconstructed pattern preserves the fringe spacing and overall contours of the data. It also features singularities (dark spots) in the already low-intensity regions of the diffraction pattern that are not clearly visible in the measured detector image. These characteristics persist across reconstructions with different random initial starting points. The apparent density of the reconstruction in 2(b) features a higher-density central region surrounded by a few-micron shell over which the density tapers off by more than an order of magnitude (Fig. S3). In an effort to understand whether the lower-magnitude outer regions seen in pale blue in 2(b) are physical or simply a reconstruction artifact, we perform a set of reconstructions with various support sizes in which the support is unaltered during the reconstruction algorithm. 2(d) compares a line cut of the data (black circles) to line cuts of the Fourier transform of the reconstructed density as a function of support diameter. The reconstructed reciprocal space intensities reproduce the features of the data well only for sufficiently large supports that extend well beyond the bounds of the higher-magnitude central region of the density. This analysis therefore suggests that there is indeed a physical basis for the lower-magnitude region featured in 2(b). Furthermore, we obtain similar results from data collected on a different SL in the same capillary (see Fig. S4, S5 and accompanying text). We defer proposing origins of this apparent lower-magnitude density region to the Discussion in order to first present the XPCS results for this same SL structure.
X-ray photon correlation spectroscopy. XPCS serves two main purposes in our study; we use it first to assess the extent of SL immobilization over the course of the measurements and, second, to characterize the nature of SL annealing. To assess the extent of SL immobilization by the quartz wool via XPCS, detector images from individual X-ray pulses were collected over a period of 140 seconds. 3(a) displays a cropped detector image of a SL (111) peak averaged over the full dataset. Following pixel calibration [40] (see Methods), the temporal autocorrelation in Fig. 3(b) was obtained from the red boxed 5 5 pixel ROI indicated in Fig. 3(a) using the multi-tau correlation algorithm, which averages shots together prior to correlation for long delays[41, 42]. We return to an analysis of the longer-delay behavior of such temporal autocorrelations below to consider the nature of the SL annealing events. At short delays, the autocorrelation in Fig. 3(b) features an oscillation, which appears as a peak in the Fourier spectrum (Fig. 3(c)) at approximately 5 Hz. Calculating an autocorrelation over each 55 pixel region on the detector and identifying its strongest-peaked frequency between 3 Hz and 10 Hz yields a frequency map of the Bragg peak in Fig. 3(d). A lower signal-to-noise ratio away from the peak makes oscillatory signals hard to resolve, and the peak frequencies in those regions vary widely. The central portion of the peak, however, consists of more homogeneous, and therefore more reliable, values with an average of Hz.
To characterize aspects of SL annealing we performed additional analysis on the dataset shown in Fig. 3. To elucidate the scale-dependent dynamics within the SL, we calculate correlation functions as a function of distance from the Bragg peak, as shown in Fig. 4(a). We divide the detector ROI containing the peak into annular regions bounded by contours of constant , and we further subdivide each region along the longitudinal and transverse components of , marked in gold and blue, respectively. In this SAXS geometry, the longitudinal direction is parallel to , and the transverse direction is tangent to the (111) diffraction ring. For each annular region, we fit the corresponding autocorrelation function, , to a compressed/stretched exponential of the form , where is a baseline, is the speckle contrast, is the decorrelation rate, and is a stretching exponent. Fig. 3(b)-(c) shows example autocorrelations and fits for respective longitudinal and transverse directions at (see also Figs. S6, S7, S8 for data and fits at additional values). Such stretched/compressed exponentials often appear in systems with heterogeneous dynamics, where the resulting distribution of time scales leads to distortion of the simple exponential decay or Gaussian correlation forms [43]. Note that a speckle contrast higher than 1 arises when there are large variations in intensity (e.g., diffraction fringes) within the ROI. Because this analysis depends on the choice of center for the annular regions of interest and there is ambiguity in how this choice should be made, we repeat the analysis for many different choices of center, using pixels within a circular region (indicated in red in Fig. 3(a)) with a 5 pixel radius around the intensity center of mass, and averaging the resultant fit parameters together. Fig. 3(d) shows the decorrelation rate as a function of for each of the radial and transverse directions, respectively. The corresponding stretching exponents, , are shown in Fig. 3(e). The behavior of at low (region i, in gray) is complex and non-monotonic. The dispersion contrasts with that of, for instance, a simple, freely-diffusing colloidal system, where [26]. Unlike a diffusive system, which has no characteristic length scale, however, our system has a clear length scale of relevance, namely the size of the SL. For values around this value we should not expect a simple dispersion. Only for sufficiently small length scales (region ii, white) do we recover standard scattering behavior of faster decorrelation at higher . In the transverse direction, the decorrelation rate is on the order of Hz and increases with increasing , while in the longitudinal direction we do not resolve such dispersive behavior (and cannot reliably extract fits at the highest values). This trend in decorrelation rate in the transverse direction, and the lack thereof in the longitudinal direction, persist even if one fits to a fixed exponent of instead (Fig. S9). In order to investigate the source of this wavevector-dependent time scale, we calculate a two-time correlation function in Fig. 3(f) using the ring-shaped region of Fig. 3(a) around Two-time analysis enables us to observe from when in the data acquisition period the time scale arises and whether it is irreversible [44, 45, 46]. The raw data were first binned down to 0.1 s intervals to more readily visualize larger jumps in intensity. We observe distinct square-like patterns in the two-time correlation, corresponding to sudden shifts in the diffraction intensity at 17 s and 90 s.
III Discussion
Having described the CDI and XPCS results, we now discuss possible interpretations and their implications for elucidating the self-assembled NCs’ products and their annealing. First, we comment on the quartz wool matrix immobilization strategy for coherent scattering characterization of individual structures in the solution phase and in particular on the sufficiency of immobilization of the structures that we interrogated. Second, we discuss the hypothesis that the discrete time scales observed in the XPCS correspond to sudden shear-like avalanches. Third, we propose and weigh various interpretations of the direct space reconstruction obtained from CDI and in particular the origin of the apparent low-density shell surrounding the higher density core. Finally, we discuss the implications from assembling our CDI and XPCS findings self-consistently by inferring the mechanism by which the observed structures are nucleated, grown, and further evolved.
This work extends the reach of coherent diffractive imaging by probing a solid, immobilized structure in its native liquid environment[47]. To be able to extract meaningful conclusions about SL structure and evolution from our diffraction data via CDI and XPCS analysis techniques, the SLs must be sufficiently immobilized in the X-ray beam. The low- behavior of the SL Bragg peak is governed by long-range order in the SL; thus the oscillations that we observe in the autocorrelation functions in 2, which feature prominently at low , are likely produced by a bulk motion of the SL, such as small librations of the SL in solution. Perturbed by solvated particles or external vibrations, the SLs sway periodically on the quartz wool, producing intensity fluctuations at a rate much slower than the X-ray pulse repetition rate of 120 Hz. While the SLs are therefore not perfectly immobilized in the capillary, these perturbations are quite small. Some challenges to extracting the SL motion from the data directly include: the time-resolved data is a 2D cross-section of the 3D diffraction pattern; rotations do not easily map onto a rectilinearly-sampled dataset, and it is not straightforward to determine intensity fluctuations from an intensity autocorrelation. We can, however, attempt to place limits on the extent of the motion using the resolution of the measured static 3D diffraction pattern. In particular, we identify sharp peaks in individual pixel intensities as a function of rocking angle, with a full-width at half-maximum of five angular steps, or 0.078 (Fig. S10). For the in-plane direction, we see fringes, normal to the transverse direction, with a full-width at half-maximum of roughly five pixels (Fig. S11). A rotation of the SL of only approximately 0.25 about the direct beam would shift a feature in the diffraction pattern by five pixels in the transverse direction. Since the data acquisition period (2 s) for each rocking angle is roughly an order of magnitude larger than the autocorrelation oscillation period, these fringe widths serve as upper bounds on the extent to which our intensity dataset may be blurred by SL librations. That is, we expect the SL rotation about the rocking axis to be no more than about 0.08 (a maximum motion of a 4 nm arc length for a 3 m structure) and the rotation about the direct beam to be no more than about 0.25 (a maximum motion of a 13 nm arc length for a 3 m structure). We cannot easily quantify rotations about a third axis orthogonal to these first two, but since there is no reason to expect that the superstructure would preferentially oscillate about such an axis, we assume a similar value for it as well. Therefore, while we cannot quantify the SL motion exactly, that we are able to resolve such sharp features in our dataset is an encouraging sign that the sample is sufficiently immobilized for the purposes of XPCS and CDI analyses.
We now turn to a discussion of the long-time behavior of the SLs, as revealed by our XPCS results. We first observe that the dispersive increase (Fig. 3(d)) at high in the transverse fit parameters extracted from the XPCS data reflects conventional dynamical behavior in which scattering decorrelates more quickly at higher wavevector () [26]. Moreover, wavevectors in the transverse direction are sensitive to shear-like motions of the SL, analogous to the sensitivity to transverse acoustic phonons near Bragg peaks[48, 49]. Therefore, since we observe this dispersion along the transverse direction but not along the longitudinal direction, we interpret the time scale as that of shear-like fluctuations of the SL. We see, furthermore, that this one-time decorrelation is accompanied by square-like patterns in the two-time correlation (Fig. 3(f)), a phenomenon that we also observe in an additional experimental dataset (Fig. S12). In a third experimental dataset (Fig. S13, S14), we do not observe any decay in the autocorrelation functions at long delay, and the corresponding two-time correlation function (Fig. S15) is highly uniform and featureless, displaying no such rapid discontinuities as those seen in Fig. 4(f). Since the absence or presence of a one-time decorrelation rate is thus accompanied by the absence or presence of these discontinuities in the two-time correlation, the decorrelation is likely dominated by sudden changes in the diffraction pattern instead of small fluctuations therein. While we cannot rule out the possibility that these sudden changes in the diffraction pattern are merely due to bulk translations or rotations of the SL in solution, the increasing dispersion in decorrelation rate at high is indicative of internal SL dynamics. Furthermore, from prior ensemble SAXS measurements on NC SLs [13], we expect SLs to continue to evolve (via growth and defect annealing) even many hours after they are initially nucleated, provided their solution phase environment is maintained. Taken together, these features in the data suggest that the high- fluctuations in the SL Bragg peak are driven by sudden shear-like “avalanches”[50, 51] of NCs rather than a continuous annealing process. Such events are roughly analogous to the microstructural avalanches previously observed in martensitic metallic alloys [46, 45]. In our previous in situ SAXS experiments of SL formation, SLs were shown to evolve at different rates depending on the quench depth at which they were formed. At shallow quenches, highly ordered SLs formed but did not evolve much after their initial formation. In contrast, at slightly deeper quenches, more disordered SLs form initially but they continue to evolve over the course of the several hour observation period[13]. This observation is consistent with a picture in which annealing proceeds via such shear-like avalanches, in which these strain-relieving events are more likely to occur in regions of high disorder[52, 53, 54]. Moreover, previous ensemble measurements also find that the SL Bragg peak width evolves according to a power law over time, a relationship we would expect to arise from the occurrence of avalanches within many SLs within an ensemble throughout a bulk measurement window[55].
Returning to the real space reconstruction obtained from CDI presented in 2 we consider interpretations of the decrease in electron density that arises moving outward from the center of the SL. One possibility is that the reconstruction algorithm has failed, assigning low non-zero values to regions of real space that extend beyond the boundary of the object. Another possibility is that the diffraction data is blurred due to the SL librations discussed above, thus yielding artifacts on large length scales in the reconstruction. A further alternative is that inhomogeneities in the beam path (due to the presence of solvent, quartz wool, and capillary glass) induce an effective decoherence in the beam and distort the diffraction signal. There are several reasons, however, to believe that the low-density regions are not an experimental artifact. First, the analysis of size-dependence presented in 2(d)suggests that there is diffracting material over a span of 2-3 m, extending beyond the high density region of the reconstruction. Additionally, we observe similar such low-density regions in the second CDI dataset that we analyzed (Figure S5). For this latter dataset the reconstructed object, including the low-density region, is smaller, in accordance with a larger fringe spacing. If beam contamination were the driving force behind a reconstruction artifact, we would expect to see its impact appear on similar length scales across all datasets. Furthermore, we can estimate an upper bound for the magnitude of phase distortion in the scattered beam. For quartz wool fibers of 9-30 m, supposing that the part of the beam passes through at worst the edge of one fiber, we would expect a relative phase shift of around . This places an upper bound on the phase distortion due to the quartz wool, and is small enough to preserve a good CDI signal (see supplemental material for quantitative estimates). One physical interpretation of these apparent lower electron densities is that they are produced by regions of greater disorder in the SL. By analyzing and comparing hypothetical ordered and disordered SL structures, we attempt to assess the impact of at least one kind of defect on such a Bragg CDI reconstruction.
Toward this end, we introduce and then thermally relax planar defects (grain boundaries) into a simulated fcc lattice, representing ordered and disordered SL structures that are 150 nm in scale. In order to reduce computational load we simulate smaller structures than in experiment. Doing so does not preclude qualitative comparison with experiment, as the grain boundaries are topological objects, and their effect on the CDI reconstruction is independent of SL size. Fig. 5(a) and 5(b) show the real space atomic positions of the original (single crystal) and polycrystalline lattices, respectively. The polycrystal’s scattering pattern is not peaked at a single reciprocal lattice vector; we choose the most intense (111) Bragg peak and plot a two-dimensional cross-section thereof in 5(d), with the (111) peak of the single crystal lattice shown in 5(c). Note that the detector image in 2(a), with its lack of centrosymmetry and its irregular fringes, more closely resembles the diffraction pattern of the polycrystalline lattice than that of the single crystal. We next take the inverse Fourier transform of a three-dimensional rectangular subregion about each diffraction peak in 5(c)-(d), using the computed scattering phases, to obtain the real space distributions shown in 5(e)-(f). While the real space density obtained from the perfect crystal is nearly constant across the object, that of the polycrystal features large variations. In particular, in domains of the polycrystal that are well-aligned with the chosen (111) peak, the density attains a value as large as that of the single crystal, whereas in misaligned regions the density drops by an order of magnitude or more. That is, in performing phase retrieval on the (111) peak alone, rather than on the full object’s Fourier transform, only those components of the object that contribute to the (111) peak appear in the reconstruction, and thus the reconstructed “electron density” is a measure of the density of planes aligned with the (111) direction at each point in the object rather than of the raw NC packing fraction. This observation also suggests that the lower-density regions are not due to residual scattering from the capillary, solvent, or quartz wool.
The combination of CDI and XPCS analysis together has proven to be a valuable means to obtain a coherent description of a self-assembled SL structure and its evolution in solution that could not otherwise have been realized. We previously found via in situ SAXS that initial SL nucleation takes place on a rapid (sub-second) time scale, and then growth and annealing progress over the course of minutes and hours. Taking all of our findings together, one interpretation of the apparent density profile we observe from CDI is that the SL crystallinity decays away from its center, with regions of greater disorder towards the surface. We would not expect such a disordered shell, however, to be so large in proportion to the crystalline core. In addition, real space optical and electron microscopy images of this system following a quench show large, micron sized polycrystalline structures (Fig. S1). Therefore a more likely explanation is that the SL (111) peak analyzed in 2 is a single crystal grain of a larger cluster of aggregated SLs, each of which nucleated separately. Our analysis of hypothetical polycrystals in 5 demonstrates that in performing CDI on a polycrystalline structure whose other grains do not meet the Bragg condition, we would expect to see regions of apparent low density in the reconstructed object. This polycrystalline picture also accords well with the account of sub-lattice dynamics indicated by correlation analysis. In particular, the presence of defects such as slip planes or grain boundaries between neighboring crystallites can lead to a build-up of strain in the SL. Once this strain becomes sufficiently large, annealing should occur via discrete, large-scale shear motions. Based on the data analyzed here and in previous work, we thus propose that Au NC SLs quickly nucleate and then coalesce into polycrystalline structures, with shear annealing occurring within and/or between crystallites via strain-relieving avalanches.
IV Conclusion
In this work, we demonstrate Bragg-CDI and XPCS on the (111) SL peak of self-assembled strongly-coupled Au NCs in the solution phase. By immobilizing the SLs in quartz wool, we are able to observe annealing-like events and to obtain SL Bragg-CDI reconstructions of their real-space density. Using these techniques we uncover a picture of SL structure and evolution that accords well with previous ensemble SAXS measurements and microscopy, in which SLs agglomerate into polycrystalline structures after initial nucleation, and become further ordered over time via shear annealing avalanches.
Only recently have such strongly-coupled SLs been synthesized from electrostatically stabilized NCs. The large parameter space at play in SL formation, if fully understood, offers unique control over the coupling between NCs and the resultant optoelectronic properties of the material, enabling access to an electronic regime beyond that of individual quantum dots or bulk semiconductors[56, 57]. The nature of the constituent NCs (e.g., metallic or semiconducting) dramatically alters the interparticle interactions, mechanism of self-assembly, and nature of the defects and annealing mechanisms within a SL. Bragg-CDI offers an incisive, in situ, real-space tool to interrogate this diverse and unexplored chemistry, which takes place on unusual length scales much larger than those of conventional materials. A deeper understanding of these facets of self-assembly will facilitate the adoption of such strongly-coupled SLs as tunable, functional materials.[58, 59, 2]
More broadly, this experiment demonstrates the successful performance of coherent diffractive imaging on a solid structure in solution. The sensitive reaction conditions at play here and in other chemical systems pose a challenge for most X-ray visualization methods, rendering it difficult to obtain detailed structural information in the structure’s native environment. By demonstrating the feasibility of CDI for in situ liquid phase measurements, this work opens the door for material imaging experiments of increasing complexity and sophistication. Furthermore, while XPCS is conventionally used for weakly-interacting systems, here we show its utility in probing supercrystalline structures on their fundamental length and time scales. Only via the combination of both XPCS and CDI together was it possible to uncover a complete picture of long-time annealing and reconstruction processes; this multimodal approach is thus a powerful means of elucidating the structural evolution of material systems at sufficiently low time scales.
V Acknowledgments
We thank Y. Cao for thoughtful discussions and advice. This work was supported by the Office of Basic Energy Sciences (BES), US Department of Energy (DOE) (award no. DE-SC0019375). Use of the LCLS is supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences under contract no. DE-AC02-76SF00515. This research also used resources of the National Synchrotron Light Source II, a U.S. Department of Energy (DOE) Office of Science User Facility operated for the DOE Office of Science by Brookhaven National Laboratory under Contract No. DE-SC0012704. C.P.N.T. was supported by the NSF (Graduate Research Fellowship no. DGE1106400). J.K.U. was supported by an Arnold O. Beckman Postdoctoral Fellowship in Chemical Sciences from the Arnold and Mabel Beckman Foundation. L.M.H. acknowledges a National Defense Science and Engineering Graduate Fellowship. A.D. was supported by a Philomathia Graduate Student Fellowship from the Kavli Nanoscience Institute at UC Berkeley. N.S.G. was supported by a David and Lucile Packard Foundation Fellowship for Science and Engineering and a Camille and Henry Dreyfus Teacher-Scholar Award.
Appendix A: Methods
V.1 SL self-assembly.
To assemble the SLs we added to a solution of colloidal gold NCs (of 4.6 nm diameter) in hydrazine in a vial in a nitrogen-filled glovebox. The solution after quenching had an Au NC concentration of 50 mg/mL and a concentration of 75 mM. Self-assembly progressed in the vial for 24 hours, after which the solution was transferred to a borosilicate glass capillary (0.9 mm outer diameter and 0.860 mm inner diameter) and sealed with inert plugs. To immobilize the SLs in their native solution, the capillary was filled with quartz wool (9-30 m in diameter) prior to the addition of the SL solution. The sealed capillaries were brought to the X-ray Correlation Spectroscopy (XCS) endstation at the Linac Coherent Light Source (LCLS) for measurement.
V.2 Time-resolved X-ray scattering.
Coherent X-ray scattering data were collected at the XCS endstation of the Linac Coherent Light Source[38, 60]. The experiment was carried out at 11.8 keV, just below the Au absorption edge, using the Si(111) monochromator in a SAXS geometry with a sample-to-detector distance of 7.7 m, measuring to 0.15 . The X-ray beam, consisting of 50 fs pulses at 120 Hz, was focused via a compound refractive lens to m and attenuated to 5 photon/s. A double-crystal Si(111) monochromator was used to produce a monochromaticity of . CDI data collection consisted of 128 angular steps over a two degree rocking range. We collected about two seconds of data collection at each angular step. All data were subject to a detector calibration pipeline consisting of gain, pedestal, threshold, and common mode corrections [40]. All shots at a given rocking angle are summed together to obtain a single image for each angular step; we then normalize each summed image by the sum of the corresponding intensity and position monitor (IPM) values, which record the intensity of the incident beam of each shot.
V.3 CDI analysis.
The Python package Phaser was used to carry out error reduction (ER), hybrid input-output (HIO), and shrinkwrap algorithms[39]. The reconstruction in 2(b) was produced using an initial support that is a sphere of radius 1.5 m. We used the following algorithm to obtain reconstructions: 3x[500 iterations ER, 200 iterations HIO], 500 iterations ER, 10x[Shrinkwrap/intersect, 5x[500 iterations ER, 200 iterations HIO], 500 iterations ER], 10x[500 iterations ER, 200 iterations HIO], 2000 iterations ER (see supplemental material for details of the Shrinkwrap/intersect step) In visualizing the reconstruction in 2(b) we use nearest-neighbor interpolation to correct for the non-unitary sampling in real space. The sampling vectors are sufficiently close to orthogonal that we do not correct for non-orthogonality. The reconstructions in 2(d) were produced in a similar manner. For more details on the reconstructions, see supplemental materials.
V.4 XPCS data analysis.
For XPCS data analysis, all one-time and two-time correlations were calculated using the Python package scikit-beam[41] using a symmetric normalization scheme. We divide the detector into ring-shaped regions of width 2.5 pixels, centered at a given point the peak, and further divide each ring into longitudinal and transverse sections. To compute the location of , we fit the (111) diffraction ring to a circle, and take the direct beam to be at the center. We calculate an autocorrelation for each ROI and fit to a stretched exponential form. We repeat this fitting process for each pixel within 5 pixels of the intensity center of mass of the time-integrated intensity pattern. We then remove outliers from the fit results (points at least 1.5 times the interquartile range larger/smaller than the upper/lower quartile) and average the remaining parameters together to get the final results.
V.5 Computational methods.
For the simulated hypothetical structures in 5, we arrange 19619 atoms into an fcc lattice. To produce the polycrystal, we randomly select 20 dividing planes that are either (111), (110), or (100) and, for each one, rotate one side of the crystal by a random angle, and shift the rotated domain a particle radius away perpendicular to the plane, to prevent particle overlap. We then minimize the energy of the resulting grain boundaries using a pairwise Lennard-Jones potential. From here, we use the NC positions to compute the structure factor and multiply by a hard-sphere form factor to obtain the complex scattering pattern. We perform a numerical inverse Fourier transform on a cubic region of interest containing the Bragg peak, the complex magnitude of which yields the density plots in Fig. 5(e) and 5(f).